OpenStudy (anonymous):

what is the integral from (x/((2x^2)+1)) as x goes to 0 to 1 dx? Could you show me step by step?

OpenStudy (anonymous):

You want to use u-substitution. So u=2x^2+1 du=4xdx (1/4)du=xdx Replacing these in your integral gives: \[\frac{1}{4}\int\limits \frac{du}{u}=\frac{1}{4}\ln|u|+C=\frac{1}{4}\ln|2x^2+1|+C\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You just need to plug your limits and get \(\frac{1}{4}[\ln(3)-\ln(1)]=\frac{1}{4}\ln(3)\).

OpenStudy (anonymous):

thank you!